Can you please provide assistance in creating the equation for this problem and also show how to solve this word problem.
Problem: Edmund is 3 years older than Frank. The sum of their ages is nineteen years. How old is each one?

Answer by Earlsdon(6294)

Let E be Edmund's age and F be Frank's age.
From the problem description, you can write:
1) E = F+3 and
2) E+F = 19
Now, you substitute the E from equation 1) into equation 2) and solve for F.
2a) (F+3) + F = 19 Simplify.
2F+3 = 19 Subtract 3 from both sides.
2F = 16 Divide both sides by 2.
F = 8 This is Frank's age.
E = F+3
E = 8+3
E = 11 This is Edmund's age.